Find the dimensions of a,band c if a is distance ,g is gravity due to acceleration and t is time s=at+BG+ct2

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
s=at+BG+ct2Dimension of at = dimension of bg dimension of at = m.sSo b(m/s2) = m.sSo dimension of B = s^3Now dimension of at = dimestion of ct^3So m.s= c.s^3So dimension of c is m/s2